OpenStudy (anonymous):

4n^2-17n+15 in fatctored form

OpenStudy (kirbykirby):

Find 2 numbers \(m\) and \(n\) such that \(m*n = 4*15 = 60\) and such that \(m+n = -17\) Then you can write your polynomial in the form: \(4n^2 + mx + nx + 15\). Then you should be able to factor 2 pairs of terms together, and then factor it again.